Rationalize the denominator
8/(√6 - 2)
I got as far as
8/(√6 - 2) * (√6 + 2)/(√6 + 2)
[8/(√6 - 2)]/(√6)^2 + (2)^2
[8/(√6 - 2)]/6+4
The book says the answer is 4√6 + 8

Answer by Cromlix(4381)

Hi there,
8/(√6 - 2)
8/(√6 - 2) * (√6 + 2)/(√6 + 2)
...........................
Note: (√6 - 2) * (√6 + 2)= 6 - 4
√6 * √6 = 6
√6 * 2 = 2√6
√6 * -2 = -2√6
2√6 + -2√6 = 0
-2 * 2 = -4
............................
8√6 + 16/6 - 4
8√6 + 16/ 2
2(4√6 + 8)/2
Cancel 2 top and bottom
= 4√6 + 8
Hope this helps :-)